# Toward Exploring the Code Understanding Capabilities of Pre-trained Code Generation Models

### v1

@@ 4.2 Q1: What is the optimal method for extracting code representations from decoder-only Transformer?
Using CodeGen (350M) as an example, we conducted an ablation experiment on the downstream task of code search, with results presented in Table 2. Our findings indicate that the optimal method for obtaining code representations varies with different padding strategies. When right padding is used, averaging all token embeddings yields better results than using only the embedding of the last token. Conversely, with left padding, the embedding of the last token outperforms the average of all embeddings. Considering results from both the CSN and CoSQA datasets, the decoder-only Transformer achieves the best representation with right padding and averaging all token embeddings. Subsequent experiments employed this setting to obtain code representations.
4.3 Q2: How well do raw pre-trained code generation models perform on code understanding tasks?
We evaluated baseline models without fine-tuning on code search and clone detection tasks. As shown in Table 4, the zero-shot performance of encoder-only models on two downstream code understanding tasks improves with the introduction of more code structure-related tasks or scaling up strategies. Notably, UniXcoder and CodeSage leverage contrastive learning tasks during pre-training, significantly enhancing zero-shot performance compared to other encoder-only models. By contrast, decoder-only models underperform on downstream tasks related to code understanding, as they primarily learn to predict the next token during the pre-training.
4.4 Q3: To what extent does contrastive learning enhance the code representations of decoder-only Transformer?
Method Code Search Clone Detection
CSN CoSQA POJ-104
Zero-shot
Encoder-only CodeBERT Feng et al. (2020) (125M) 0.10 0.24 20.38
GraphCodeBERT Guo et al. (2021) (125M) 11.26 16.20 31.27
UniXcoder Guo et al. (2022) (125M) 46.40 42.11 42.08
CodeSage Zhang et al. (2024) (1.3B) 71.24 47.53 73.07
Decoder-only CodeGPT Lu et al. (2021) (125M) 0.12 0.04 9.41
CodeGen Nijkamp et al. (2023) (350M) 1.42 0.45 13.20
SantaCoder Allal et al. (2023) (1.1B) 0.07 0.11 15.57
phi-1 Gunasekar et al. (2023) (1.3B) 6.66 6.75 27.62
[…]
CodeGen Nijkamp et al. (2023) (350M) + CL4D 72.00 ($\uparrow$ 70.58) 51.2 ($\uparrow$ 50.75) 45.84 ($\uparrow$ 32.64)
SantaCoder Allal et al. (2023) (1.1B) + CL4D 74.18 ($\uparrow$ 74.11) 52.8 ($\uparrow$ 52.69) 71.14 ($\uparrow$ 55.57)
phi-1 Gunasekar et al. (2023) (1.3B) + CL4D 73.20 ($\uparrow$ 66.54) 52.9 ($\uparrow$ 46.15) 69.99 ($\uparrow$ 42.37)
Fine-tune
Encoder-only CodeBERT Feng et al. (2020) (125M) 70.17 65.7 83.79
GraphCodeBERT Guo et al. (2021) (125M) 72.10 68.4 85.50
UniXcoder Guo et al. (2022) (125M) 74.40 70.1 89.56
CodeSage Zhang et al. (2024) (1.3B) 75.80 68.0 87.70
Decoder-only CodeGPT Lu et al. (2021) (125M) + CL4D 70.20 69.0 87.96
CodeGen Nijkamp et al. (2023) (350M) + CL4D 73.30 71.5 89.68
SantaCoder Allal et al. (2023) (1.1B) + CL4D 74.98 72.2 83.98
phi-1 Gunasekar et al. (2023) (1.3B) + CL4D 75.83 72.8 92.72
Table 4: Comparison of the performance of encoder-only models and decoder-only models on code understanding tasks in zero-shot and fine-tune settings. Bold indicates optimal performance, and bold and underlined indicates suboptimal performance.
Model Python Java Go Php JavaScript Ruby Avg
Encoder-only CodeBERT Feng et al. (2020) (125M) 0.05 0.03 0.14 0.02 0.04 0.34 0.10
GraphCodeBERT Guo et al. (2021) (125M) 10.37 8.59 12.47 8.07 7.29 20.79 11.26
UniXcoder Guo et al. (2022) (125M) 42.17 43.92 61.39 35.21 40.46 55.22 46.40
CodeSage Zhang et al. (2024) (1.3B) 70.77 70.21 83.71 61.33 69.50 71.92 71.24
Decoder-only CodeGPT Lu et al. (2021) (125M) + CL4D 67.1 66.3 83.1 58.0 61.8 69.1 67.57
CodeGen Nijkamp et al. (2023) (350M) + CL4D 70.6 71.5 86.5 63.0 67.3 72.9 72.00
SantaCoder Allal et al. (2023) (1.1B) + CL4D 73.3 72.9 86.6 65.3 71.4 75.6 74.18
phi-1 Gunasekar et al. (2023) (1.3B) + CL4D 72.9 71.6 86.3 63.8 69.1 75.5 73.20
Table 5: Results of code search task over six programming languages of CSN in the zero-shot setting.
The experimental results from Q2 indicate that the extensive code knowledge acquired by pre-trained code understanding models remains underutilized. To leverage this knowledge for code understanding tasks, we applied the CL4D method described in Section 3 for continued pre-training of these decoder-only models. The experimental outcomes, presented in Table 4, demonstrate that: 1) Our proposed method significantly enhances the zero-shot performance of decoder-only models on both code search and clone detection tasks, with average improvements ranging from 40% to 60%, and a maximum improvement of 74.11%. 2) Our method enables the zero-shot and fine-tune performance of decoder-only models on various code understanding tasks to surpass that of encoder-only models of comparable size. 3) Our method enables decoder-only models to achieve performance comparable to state-of-the-art, fine-tuned encoder-only models of the same size on the CSN dataset, even without fine-tuning. 4) As the number of parameters in decoder-only models increases, their performance on downstream code understanding tasks progressively improves.
4.5 Q4: Why does our approach work?
The experimental results from Q3 indicate that our proposed method significantly enhances the performance of existing decoder-only models on various downstream code understanding tasks, both in zero-shot and fine-tuning scenarios. To investigate the effectiveness of our method, we designed ablation experiments to analyze the contributions of different components in CL4D. Additionally, we visualized the impact of our method on the semantic encoding capabilities of decoder-only models to provide a more intuitive understanding of its effectiveness.
@@ 4.5 Q4: Why does our approach work?
Ablation Study We trained three groups of models based on the CodeGen model: 1) the original Codegen; 2) the model trained using the contrastive learning method with in-batch negatives; 3) the model trained using the contrastive learning method with in-batch negatives and hard negatives. We compared their zero-shot performance on downstream code understanding tasks, with results shown in Table 6. The findings reveal that using in-batch negatives significantly enhances the code understanding ability of decoder-only models. Adding hard negatives further improves performance by approximately 1.5%. Contrastive learning is expected to become an effective method for the decoder-only Transformer architecture to unify code understanding and code generation tasks.
@@
Table 6: Ablation study results of CL4D in zero-shot setting.
@@
Visualization To explore why CL4D effectively improves the performance of decoder-only models on code understanding tasks, we visualized the representation distribution of CodeGPT and CodeGPT+CL4D on CSN, CoSQA, and POJ-104 datasets. For the CSN dataset, each query has only one correct answer, so we randomly sampled 30 query-code pairs from the test set to form 30 categories, with 2 samples in each category. The same approach was applied to the CoSQA dataset. In the POJ-104 test set, which has 24 problems, each with multiple correct answers, we sampled 5 correct answers per problem, forming 24 categories with 5 samples each.
We then feed the sampled data into the models to obtain their high-dimensional feature space representations. Using t-SNE Van der Maaten and Hinton (2008), we reduce the dimensionality of these representations and visualize their distribution, as shown in Figure 4. Points with the same color and shape represent samples with the same semantics. Ideally, their representations in the vector space should be as close as possible. The results show that the representations encoded by the original decoder-only model are highly scattered, with semantically similar samples far apart in the vector space. The decoder-only model pre-trained with CL4D aligns the representations of semantically similar samples, bringing them closer together in the semantic space.
@@
5 Conclusion
In this paper, we have presented a novel application of pre-trained decoder-only code generation models to code understanding tasks. We explore four distinct methods by which these decoder-only models develop code representations. Our study includes a comparative analysis of encoder-only and decoder-only models, examining their performance on downstream code understanding tasks in a zero-shot setting. To enhance the representational capabilities of the decoder-only model, we introduce a contrastive learning method, CL4D, which constructs in-batch negatives and hard negatives for training. Experimental results demonstrate that CL4D enables decoder-only models to significantly outperform encoder-only models in most tasks. Our findings indicate that decoder-only Transformer have the potential to unify code understanding and generation tasks. We will focus on further exploring this direction to preserve the code generation capabilities of larger decoder-only models while enhancing their code understanding abilities in the future work.

### v2

@@ Introduction
Using CodeGen (350M) as an example, we conducted an ablation experiment on the downstream task of code search, with results presented in Table 2. Our findings indicate that the optimal method for obtaining code representations varies with different padding strategies. When right padding is used, averaging all token embeddings yields better results than using only the embedding of the last token. Conversely, with left padding, the embedding of the last token outperforms the average of all embeddings. Considering results from both the CSN and CoSQA datasets, the decoder-only Transformer achieves the best representation with right padding and averaging all token embeddings. This setting was used in all other experiments.
RQ3: Ablation study
To investigate the effectiveness of our method, we designed ablation experiments to analyze the contributions of different components in CL4D.
@@
We trained three groups of models based on the CodeGen model: 1) the original Codegen; 2) the model trained using the contrastive learning method with in-batch negatives; 3) the model trained using the contrastive learning method with in-batch negatives and hard negatives. We compared their zero-shot performance on downstream code understanding tasks, with results shown in Table 3. The findings reveal that using in-batch negatives significantly enhances the code understanding ability of decoder-only models. Adding hard negatives further improves performance by approximately 1.5%. Contrastive learning is expected to become an effective method for the decoder-only Transformer architecture to unify code understanding and code generation tasks.
@@
Table 3: Ablation results of CL4D in the zero-shot setting.
RQ4: Zero-shot
We evaluated baseline models without fine-tuning on code search and clone detection tasks. As shown in Table 4, the zero-shot performance of encoder-only models on two downstream code understanding tasks improves with the introduction of more code structure-related tasks or scaling up strategies. Notably, UniXcoder and CodeSage leverage contrastive learning tasks during pre-training, significantly enhancing zero-shot performance compared to other encoder-only models. By contrast, decoder-only models underperform on downstream tasks related to code understanding, as their pre-training primarily focuses on next-token prediction rather than bidirectional semantic representation.
The experimental results indicate that the extensive code knowledge acquired by pre-trained code understanding models remains underutilized. To leverage this knowledge for code understanding tasks, we applied the CL4D method described in Section Methodology for continued pre-training of these decoder-only models. The experimental outcomes, presented in Table 4, demonstrate that: 1) Our proposed method significantly enhances the zero-shot performance of decoder-only models on both code search and clone detection tasks, with average improvements ranging from 40% to 60%, and a maximum improvement of 75.90%. 2) Our method enables the zero-shot performance of decoder-only models on various code understanding tasks to surpass that of encoder-only models of comparable size. 3) Our method enables decoder-only models to match the performance of same-sized, fine-tuned encoder-only state-of-the-art models on the CSN dataset, even without fine-tuning.
Method Code Search Clone Detection
CSN CoSQA POJ-104
Encoder-only CodeBERT (Feng et al. 2020) (125M) 0.10 0.24 20.38
GraphCodeBERT (Guo et al. 2021) (125M) 11.26 16.20 31.27
UniXcoder (Guo et al. 2022) (125M) 46.40 42.11 42.08
CodeSage (Zhang et al. 2024) (1.3B) 71.24 47.53 73.07
Decoder-only CodeGPT (Lu et al. 2021) (125M) 0.12 0.04 9.41
CodeGen (Nijkamp et al. 2023) (350M) 1.42 0.45 13.20
SantaCoder (Allal et al. 2023) (1.1B) 0.07 0.11 15.57
phi-1 (Gunasekar et al. 2023) (1.3B) 6.66 6.75 27.62
DeepSeek-Coder (Guo et al. 2024) (1.3B) 0.12 0.63 16.51
Decoder-only CodeGPT (Lu et al. 2021) (125M) + CL4D 67.56 ($\uparrow$ 67.44) 53.49 ($\uparrow$ 53.45) 25.93 ($\uparrow$ 16.52)
CodeGen (Nijkamp et al. 2023) (350M) + CL4D 71.97 ($\uparrow$ 70.55) 51.18 ($\uparrow$ 50.73) 45.84 ($\uparrow$ 32.64)
SantaCoder (Allal et al. 2023) (1.1B) + CL4D 74.18 ($\uparrow$ 74.11) 52.82 ($\uparrow$ 52.71) 71.14 ($\uparrow$ 55.57)
phi-1 (Gunasekar et al. 2023) (1.3B) + CL4D 73.21 ($\uparrow$ 66.55) 52.95 ($\uparrow$ 46.20) 69.99 ($\uparrow$ 42.37)
DeepSeek-Coder (Guo et al. 2024) (1.3B) + CL4D 76.02 ($\uparrow$ 75.90) 48.34 ($\uparrow$ 47.71) 71.18 ($\uparrow$ 54.67)
Table 4: Comparison of the performance of encoder-only models and decoder-only models on code understanding tasks in the zero-shot settings. Bold indicates optimal performance, and bold and underlined indicates suboptimal performance.
RQ5: Why does our approach work?
Figure 3: Case study on the code search task in the zero-shot setting.
Case study Figure 3 displays the top-1 ranked code snippets returned by DeepSeek-Coder, CodeSage, and DeepSeek-Coder+CL4D for the query “python only list files with specific extension”. DeepSeek-Coder+CL4D successfully returns the correct code snippet, which iterates through files and identifies those matching the specified extension. In contrast, the code snippets from DeepSeek-Coder and CodeSage are incorrect: DeepSeek-Coder returns a result entirely unrelated to the query, while CodeSage’s snippet only identifies the extension without returning the corresponding files. These results indicate that continued pre-training of decoder-only models with CL4D significantly enhances their code understanding ability, allowing them to outperform encoder-only models of the same size.
@@
Visualization To investigate why CL4D effectively improves the performance of decoder-only models on code understanding tasks, we visualized the representation distribution of CodeGPT and CodeGPT+CL4D across the CSN, CoSQA, and POJ-104 datasets. For the CSN dataset, each query has only one correct answer, we randomly sampled 30 query-code pairs from the test set to form 30 categories, each containing 2 samples. The same approach was applied to the CoSQA dataset. For the POJ-104 test set, which has 24 problems, each with multiple correct answers, we sampled 5 correct answers per problem, forming 24 categories with 5 samples each.
We then feed the sampled data into the models to obtain their high-dimensional feature space representations. Using t-SNE (Van der Maaten and Hinton 2008), we reduce the dimensionality of these representations and visualize their distribution, as shown in Figure 4. Points with the same color and shape represent samples with the same semantics. Ideally, their representations in the vector space should be as close as possible. The results show that the representations encoded by the original decoder-only model are highly scattered, with semantically similar samples far apart in the vector space. The decoder-only model pre-trained with CL4D aligns the representations of semantically similar samples, bringing them closer together in the semantic space.
@@
Related works
Code Represention Recent work has focused on obtaining programming language representations for code understanding tasks. Models like CodeBERT (Feng et al. 2020), StarEncoder (Li et al. 2023), CodeSage (Zhang et al. 2024), and CoLSBERT (Lin et al. 2024) use the Masked Language Modeling (MLM) objective from NLP, applying it to serialized code. Recognizing the structural differences between code and text, some models enhance learning by incorporating code structures. For example, GraphCodeBERT (Guo et al. 2021) uses data flow, TreeBERT (Jiang et al. 2021) utilizes the abstract syntax tree (AST), and SynCoBERT (Wang et al. 2021) and UniXcoder (Guo et al. 2022) include serialized ASTs during training.
Although these models advance code understanding, they remain relatively small, typically trained on datasets with only millions of samples, and even the largest model, CoLSBERT, has only 1.5B parameters. While scaling up model size and training data can further improve performance, it also incurs substantial computational and resource costs. Inspired by the success of large-scale decoder-only code generation models, our work aims to leverage such pre-trained models to improve code understanding tasks, achieving state-of-the-art performance.
Contrastive Learning for Code Contrastive learning has been widely employed in representation learning for deep learning networks. Recent studies have extended this paradigm to code corpora, proposing various methods to construct positive or negative samples to enhance code representation. For example, CoSQA (Huang et al. 2021) constructs positive samples by rewriting the query, SynCoBERT (Wang et al. 2021) and Code-MVP (Wang et al. 2022) create positive sample pairs across different code modalities using diverse information obtained during program compilation, and UniXcoder (Guo et al. 2022), inspired by SimCLR (Chen et al. 2020), uses dropout to construct positive samples in the latent vector space. Additionally, ChatDANCE (Wang et al. 2023) uses large language models such as ChatGPT to synthesize positive code samples. CodeRetriever (Li et al. 2022), R2 (Dong et al. 2023), and CodeSage (Zhang et al. 2024) construct hard negatives for queries using rankers. In this paper, to mitigate the limited representation capability of decoder-only models by their unidirectional attention mechanism, we select code snippets with different semantics from the query but close in the representation space as hard negatives.
Conclusion
In this paper, we have presented a novel application of pre-trained decoder-only code generation models to code understanding tasks. We explore four distinct methods by which these decoder-only models develop code representations. Our study includes a comparative analysis of encoder-only and decoder-only models, examining their performance on downstream code understanding tasks in a zero-shot setting. To enhance the representational capabilities of the decoder-only model, we introduce a contrastive learning method, CL4D, which constructs in-batch negatives and hard negatives for training. Experimental results demonstrate that CL4D enables decoder-only models to significantly outperform encoder-only models in most tasks. Our findings indicate that decoder-only Transformer have the potential to unify code understanding and generation tasks.